OpenStudy (anonymous):

Simplify: (n^4)3/2 Question 20 options: 6 n7 n5 n6

OpenStudy (mathmate):

Hint: \(\huge (a^b)^\frac{c}{d}=a^{\frac{bc}{d}}\)